# Relativistic heavy-ion collisions; classical Yang-Mills theory; glasma

### 09 Jan 2024 v1

@@ 2.2 $3$D glasma in collision of relativistic nuclei
where $(A_{\tilde{\tau}},A_{\tilde{\eta}})$ are related to $(A^{-},A^{+})$ through the general coordinates trasformation. This solution satisfies the Fock-Schwinger gauge condition for the modified proper time, $A_{\tilde{\tau}}=0$. Accordingly, the electric fields in this modified Fock-Schwinger gauge condition are given by
$E^{i}$ $=\frac{1}{\tilde{\tau}}\partial_{\tau}A_{i}=x^{-}\partial_{-}A^{(1)}_{i}+x^{+}\partial_{+}A^{(2)}_{i}\ ,$ (20)
$E^{\tilde{\eta}}$ $=\tilde{\tau}\partial_{\tau}A_{\tilde{\eta}}=0\ .$ (21)
However, in the actual setup for calculations performed later, we assume that the color charge densities before the collision have a Gaussian shape. Since the Gaussian function extends to infinity, two nuclei are correlated with each other even in the infinite past. Therefore, the simple sum of the solutions for each single nucleus, as described above, is no longer a solution of the evolution equations and violates Gauss’s law,
@@
The additional longitudinal electric field is negligibly small, $E^{\tilde{\eta}}\ll E^{i}$ because the overlap of the tails of the Gaussian-shaped color charge densities is very minor at $\tilde{\tau}=\tilde{\tau}_{\rm ini}$. Finally, the initial condition of the classical color charge density and CYM field we employ is given by the sum of the solutions of the evolution equations for each individual nucleus with the small modification for the longitudinall electric field,
@@
We assume the classical color charge density in the infinite past as the initial condition of the color charge density $\rho_{\rm ini}$, under the premise that each nucleus is still largely unaffected by the other at the initial simulation time $\tilde{\tau}_{\rm ini}$. To simulate the effect of non-trivial longitudinal correlation in the classical color charge density, the initial condition $\rho_{\rm ini}$ is modeled using the modified MV model. In this model, $\rho_{\rm ini}$ is given to satisfy the following relation, the generalization of Eq. (8) by replacing the delta function for $x^{\mp}$ with a Gaussian function,
@@
where $N(x^{\mp},l^{(1/2)}_{\rm L})$ denotes the normal distribution with a variance $l^{(1/2)}_{\rm L}$, and $g^{2}\mu^{(1/2)}$ becomes a function of the center position of the two color charge densities. The variances of the Gaussian functions, $l^{(1/2)}_{\rm L}$, act as longitudinal correlation lengths. For simplicity, we further assume that the averaged square of the color charge density also has a Gaussian shape.
@@
Under the setup shown in Eq. (28), the event-by-event color charge density satisfying the relation shown in Eqs. (27) as an event average can be obtained by using the event-by-event random number $\Gamma$,
@@
where $k_{\perp}$ is the transverse momentum. The ultraviolet cutoff $\Lambda_{\rm UV}$ is also introduced to facilitate taking a continuum limit. The ultraviolet cutoff should be chosen as a larger value than the saturation scale $Q_{\rm s}$, which is the scale focused on in the CGC.
@@ 3 Numerical Results
We show the numerical results for the 3D glasma, computed using the modified Milne coordinates. Our simulations are conducted using a lattice framework. We utilize the leap-frog method to solve the discretized evolution equations set on the lattice. As for the boundary conditions, we employ periodic boundaries in the transverse directions, while ensuring that both the CYM field and classical color current are set to zero at the longitudinal direction’s boundaries.
In what follows, we first present a set of parameters involved in the model. Then, we provide the numerical results for the pressure and energy density at mid-rapidity, as well as the pressure results in the broader rapidity region. The energy density, transverse pressure and longitudinal pressure are defined by the energy-momentum tensor in the ’usual’ Milne coordinates,
@@
Here they are integrated values over the transverse plane since we focus on their dependence on rapidity. Reflecting the fact that the energy-momentum tensor of the CYM field is traceless due to its conformal symmetry, the relationship between energy and pressure, $\varepsilon=2P_{\perp}+P_{\rm L}$, is held. Although our actual simulations are conducted in the modified Milne coordinates, we can obtain results in the usual Milne coordinates through a general coordinate transformation of the energy-momentum tensor,
$T^{\tau\tau}$ $=\frac{1}{\tilde{\tau}^{2}-2\tilde{\tau}\bar{x}_{\rm c}\cosh{\tilde{\eta}}+\bar{x}^{2}_{\rm c}}\Big{[}\left(\tilde{\tau}-\bar{x}_{\rm c}\cosh{\tilde{\eta}}\right)^{2}T^{\tilde{\tau}\tilde{\tau}}$
$+\left(\bar{x}_{\rm c}\sinh{\tilde{\eta}}\right)^{2}\tilde{\tau}^{2}T^{\tilde{\eta}\tilde{\eta}}-\bar{x}_{\rm c}\sinh{\tilde{\eta}}\left(\tilde{\tau}-\bar{x}_{\rm c}\cosh{\tilde{\eta}}\right)\tilde{\tau}T^{\tilde{\tau}\tilde{\eta}}\Big{]}\ ,$ (36)
$\tau^{2}T^{\eta\eta}$ $=\frac{1}{\tilde{\tau}^{2}-2\tilde{\tau}\bar{x}_{\rm c}\cosh{\tilde{\eta}}+\bar{x}^{2}_{\rm c}}\Big{[}\left(\bar{x}_{\rm c}\sinh{\tilde{\eta}}\right)^{2}T^{\tilde{\tau}\tilde{\tau}}$
$+\left(\tilde{\tau}-\bar{x}_{\rm c}\cosh{\tilde{\eta}}\right)^{2}\tilde{\tau}^{2}T^{\tilde{\eta}\tilde{\eta}}-\bar{x}_{\rm c}\sinh{\tilde{\eta}}\left(\tilde{\tau}-\bar{x}_{\rm c}\cosh{\tilde{\eta}}\right)\tilde{\tau}T^{\tilde{\tau}\tilde{\eta}}\Big{]}\ ,$ (37)
where $\bar{x}_{\rm c}=\sqrt{2}x_{\rm c}$.
3.1 Model paramteres
@@
The set of model parameters for the initial $3$D color charge density includes: the longitudinal extension of the classical color charge density $r_{L}$, the longitudinal correlation length in the classical color charge density $l_{L}$, the parameter controlling the averaged square of the color charge density $g^{2}\bar{\mu}$, and the infrared and ultraviolet cutoffs, $\Lambda_{\rm IR}$ and $\Lambda_{\rm UV}$. To simulate the Au-Au collision with a center-of-mass energy per nucleon pair of $200$ GeV at RHIC, we adjust $\sqrt{2}r_{L}$ to roughly match the radius of a gold nucleus divided by the gamma factor $\gamma=108$ and the factor $\sqrt{2}$ stemming from the definition of the light-cone variables, $r_{L}=0.1/Q_{\rm s}$, where the saturation scale is set at $Q_{\rm s}=1$ GeV. The parameter $g^{2}\bar{\mu}$ is set to ensure the integration of the averaged square of the color charge density over the light-cone variable coincides with the squared saturation scale, $\int z\left(g^{2}\mu(x,\bm{x}_{\perp})\right)^{2}=Q^{2}_{s}$. This setup for $g^{2}\bar{\mu}$ indicates the event average of a given observable has a translational invariance in the transverse plane. Such a translational invariance is approximately established in the central region of large nuclei in the transverse plane, where the thickness of the nucleus does not vary largely. Therefore, our setup simulates the glasma in the central region, as depicted in Fig. 1. The infrared cutoff is set at $\Lambda_{\rm IR}=0.2Q_{\rm s}$, and the ultraviolet cutoff is taken to be double the value of $Q_{\rm s}$, $\Lambda_{\rm UV}=2Q_{\rm s}$. It is important to note that in the classical approximation, the choice of the coupling constant only provides the trivial change in the field variables, as $A,E,J\propto 1/g$. This is because the coupling constant can be excluded in both the classical equations of motion and the continuity equation by redefining the scaled fields, $A^{\prime}=gA,E^{\prime}=gE$ and $J^{\prime}=gJ$. The numerical results shown in the later sections are normalized by $g$ to make them independent of $g$.
We explore the effect of longitudinal correlation by comparing the 3+1D glasma under various correlation lengths $l_{\rm L}$, ranging from $0.01r_{\rm L}$ to $2r_{\rm L}$. The smallest value $l_{\rm L}=0.01r_{\rm L}$, significantly less than the nucleus radius, could be thought to arise from the intricate structure within a nucleon. The largest value $l_{\rm L}=2r_{\rm L}$, twice the nucleus radius, is the maximum limit discussed in Sec. 2.3. It is crucial to emphasize that the physical origin of the longitudinal correlation is beyond the scope of this paper.
Also, the following parameters should be inputted for the 3D glasma simulation. These parameters are chosen to be sufficiently safe values to minimize their impact on the physical results. The transverse lattice grid number and lattice spacing are set to $N_{\perp}=56$ and $a_{\perp}=0.4/Q_{s}$, respectively. For the lattice in the modified rapidity direction $\tilde{\eta}$, the grid number and spacing are $L_{\tilde{\eta}}=896,7168$, and $a_{\tilde{\eta}}=2.6/L_{\tilde{\eta}}$, respectively. It is noteworthy that simulations with smaller $l_{\rm L}$ require smaller longitudinal lattice spacings to accurately capture the structure of the color charge density in the longitudinal direction. The beginning modified proper time in the simulation and the center positions of the nuclei are taken as $\tilde{\tau}_{\rm ini}=32r_{\rm L}$ and $x_{\rm c}=20\sqrt{2}r_{\rm L}$, respectively. All results in this section are estimated using $48$ events. The error in the results is estimated by dividing the unbiased variance by the square root of the number of events. According to the central limit theorem, the unbiased variance aligns with the variance of the mean value in the infinitely large number of events.
@@ 3.2 Midrapidity region
First, we investigate the behavior of pressure and energy density in the mid-rapidity region, $\eta=0$. The figure 2 illustrates the evolution of glasma pressure at $\eta=0$, with the pressure normalized to the energy density. The time range, $0\leq Q_{\rm s}\tau\leq 9.7$, covers the time before the onset time of the hydrodynamics, $\tau=0.6\sim 1.0$ Heinz , predicted by the experimental analysis. The blue and red lines indicate pressures resulting from collisions of nuclei with two different lengths of longitudinal correlation in color charge densities, $l_{\rm L}=0.01r_{\rm L}$ and $2r_{\rm L}$ The solid and dotted lines correspond to transverse and longitudinal pressures over the energy density, $P_{\perp}/\varepsilon$ and $P_{\rm L}/\varepsilon$, respectively.
Despite exploring two distinct longitudinal correlations in the color charge densities, the observed pressures were qualitatively similar, though they showed quantitative differences, particularly in the early time region. Notably, at $Q_{\rm s}\tau=0$, transverse pressure is minimal compared to longitudinal pressure, reflecting the overlap of two nuclei centers at $(\tau,\eta)=(0,0)$ and the dominance of the Weizsäcker-Williams field, a transverse electromagnetic field representing of nuclei, which significantly contributes to longitudinal but not transverse pressure. Between $Q_{\rm s}\tau=0$ and $0.6$, both transverse and longitudinal pressures undergo substantial changes, eventually acquiring opposite signs. The realization of the relation, $P_{\perp}>0>P_{\rm L}$, indicates the formation of a strong longitudinal color electromagnetic field between separating nuclei, like the boost invariant glasma. From $Q_{\rm s}\tau=0.6$ to $2$, the transverse pressure trends towards $0.5$, while longitudinal pressure approaches zero, with no significant changes observed beyond $\tau=2$. These findings are in qualitative agreement with the previous results from the glasma calculations, where other effects of boost invariance breaking are taken into account Instability14 ; Instability15 ; Instability16 ; Instability17 ; Instability18 ; Instability19 ; Instability20 ; Instability21 ; 3DJIMWLK .
@@
The figure 3 illustrates the evolution of the glasma energy density $\varepsilon$ at $\eta=0$ normalized by that for $Q_{\rm s}\tau=1$, $\varepsilon_{0}$, where the solid blue and red lines depict the energy density arising from collisions of nuclei with two different lengths of longitudinal correlation in color charge densities, $l_{\rm L}=0.01r_{\rm L}$ and $2r_{\rm L}$. The time range is taken to be the same as in Fig. 3. As same as the pressure isotropizatoion in Fig. 3, despite of the choice of $l_{\rm L}$, the results of the $3D$ glasma simulation are same qualitatively at late time. Between $Q_{\rm s}\tau=0$ and $0.6$, the energy density is dominated by the Weizsäcker-Williams field rather than the glasma. After $Q_{\rm s}\tau=0.6$, as the nuclei move past $\eta=0$, the evolution of the glasma’s energy density becomes apparent. A power function fitting with fitting parameters $A$ and $B$, $f_{\rm fit}(\tau)=A/\tau^{B}$, was employed to accurately determine the decay rate of the glasma energy density in the later stages. The obtained fitting parameter $B$, $B=1.06$ for $l_{\rm L}=0.01r_{\rm L}$ and $B=1.04$ for $l_{\rm L}=2r_{\rm L}$, indicates that the energy of the glasma at $\eta=0$ decrease as almost the inverse of the proper time, $1/\tau$. This $1/\tau$ decrease in the energy density at late time has been observed by previous $2D$ and $3D$ glasma calculations Instability14 ; Instability16 ; Instability22 .
The negligible longitudinal pressure, $P_{\perp}\gg P_{\rm L}$, and the decay of energy density, $\epsilon\propto 1/\tau$, in the late stage, as observed above, are consistent with the system expanding longitudinally like a dilute gas. These observations significantly deviate from the behavior expected in an ideal gas, where pressure is isotropic $P_{\perp}=P_{\rm L}=P$, and the energy density decay is predicted by Bjorken hydrodynamics using the ideal gas equation of state, $\varepsilon=3P$, $\varepsilon\propto 1/\tau^{4/3}$.

### 29 Feb 2024 v3

@@ 2.2 $3$D Glasma in Collision of Relativistic Nuclei
where $(A_{\tilde{\tau}},A_{\tilde{\eta}})$ are related to $(A^{-},A^{+})$ through the general coordinate transformation. This solution satisfies the Fock–Schwinger gauge condition for the modified proper time, $A_{\tilde{\tau}}=0$. Accordingly, the electric fields in this modified Fock–Schwinger gauge condition are given by
$E^{i}$ $=\tilde{\tau}\partial_{\tilde{\tau}}A_{i}=x^{-}\partial_{-}A^{(1)}_{i}+x^{+}\partial_{+}A^{(2)}_{i}\ ,$ (20)
$E^{\tilde{\eta}}$ $=\frac{1}{\tilde{\tau}}\partial_{\tilde{\tau}}A_{\tilde{\eta}}=0\ .$ (21)
However, in the actual setup for calculations performed later, we assume that the color charge densities before the collision have a Gaussian shape. Since the Gaussian function extends to infinity, two nuclei are correlated with each other, even in the infinite past. Therefore, the simple sum of the solutions for each single nucleus, as described above, is no longer a solution of the evolution equations and violates Gauss’s law,
@@
The additional longitudinal electric field is negligibly small, $E^{\tilde{\eta}}\ll E^{i}$, because the overlap of the tails of the Gaussian-shaped color charge densities is very minor at $\tilde{\tau}=\tilde{\tau}_{\rm ini}$. Finally, the initial condition of the classical color charge density and CYM field we employ is given by the sum of the solutions of the evolution equations for each individual nucleus with the small modification for the longitudinall electric field,
@@
We assume the classical color charge density in the infinite past as the initial condition of the color charge density, $\rho_{\rm ini}$, under the premise that each nucleus is still largely unaffected by the other at the initial simulation time, $\tilde{\tau}_{\rm ini}$. To simulate the effect of non-trivial longitudinal correlation in the classical color charge density, the initial condition, $\rho_{\rm ini}$, is modeled using the modified MV model. In this model, $\rho_{\rm ini}$ is given to satisfy the following relation, the generalization of Equation (8) by replacing the delta function for $x^{\mp}$ with a Gaussian function,
@@
where $N(x^{\mp},l^{(1/2)}_{\rm L})$ denotes the normal distribution with a variance of $l^{(1/2)}_{\rm L}$, which represents the longitudinal shape of nucleus in the $x^{\mp}$ coordinate, and $g^{2}\mu^{(1/2)}$ becomes a function of the center position of the two color charge densities. The variances of the Gaussian functions, $l^{(1/2)}_{\rm L}$, act as longitudinal correlation lengths. For simplicity, we further assume that the averaged square of the color charge density also has a Gaussian shape,
@@
Under the setup shown in Equation (28), the event-by-event color charge density satisfying the relation shown in Equation (27) as an event average can be obtained by using the event-by-event random number $\Gamma$,
@@
where $k_{\perp}$ is the transverse momentum. The ultraviolet cutoff, $\Lambda_{\rm UV}$, is also introduced. The UV cutoff suppresses the high transverse momentum modes of the color charge density that exceed the limit of the CGC description. The color charge density represents the aggregate of color charges from hard partons within a transverse resolution scale, which is approximately the inverse of the transverse momentum. Therefore, UV modes originating from color charges in a small transverse spatial region are not so dense that the classical treatment is justifiable. Furthermore, the UV cutoff plays a role in facilitating the taking of the limit in continuous space. It is shown by an analytic calculation using the MV model that, without UV cutoff, a local operators of two gauge fields diverge in continuous space CFKL . The ultraviolet cutoff should be chosen as a larger value than the saturation scale, $Q_{\rm s}$, which is the scale focused on in the CGC. A detailed analysis of the UV cutoff dependence is beyond the scope of this paper.
@@ 3 Numerical Results
We show the numerical results for the 3D glasma, computed using the modified Milne coordinates. Our simulations are conducted using a lattice framework with a link variable, which is a gauge covariant and ensures that Gauss’s law holds under their evolution equations 3Dglasma6 . We utilize the leap-frog method to solve those equations set on the lattice. As for the boundary conditions, we employ periodic boundaries in the transverse directions, while ensuring that both the CYM field and classical color current are set to zero at the longitudinal direction’s boundaries. The number of colors is $2$, not $3$. While calculations based on SU($2$) cannot be used to make quantitative predictions for relativistic heavy-ion experiments, they are expected to lead to results qualitatively similar to SU($3$).
In what follows, we first present the set of parameters involved in the model. Then, we provide the numerical results for the pressure and energy density at mid-rapidity, as well as the pressure results in the broader rapidity region. The energy density, transverse pressure, and longitudinal pressure are defined by the energy–momentum tensor in the ’usual’ Milne coordinates,
@@
Here, they are integrated values over the transverse plane since we focus on their dependence on rapidity. Reflecting the fact that the energy–momentum tensor of the CYM field is traceless due to its conformal symmetry, the relationship between energy and pressure, $\varepsilon=2P_{\perp}+P_{\rm L}$, is held. Since the actual simulation is performed in modified Milne coordinates, the energy–momentum tensor to be calculated directly is in modified Milne coordinates $T^{\tilde{\tau}\tilde{\tau}},T^{\tilde{\tau}\tilde{\eta}}$ and $T^{\tilde{\eta}\tilde{\eta}}$. Therefore, the energy–momentum tensor in usual Milne coordinates $T^{\tau\tau}$ and $T^{\eta\eta}$ can be obtained by a general coordinate transformation,
$\begin{split}T^{\tau\tau}&=\frac{1}{\tilde{\tau}^{2}-2\tilde{\tau}\bar{x}_{\rm c}\cosh{\tilde{\eta}}+\bar{x}^{2}_{\rm c}}\Big{[}\left(\tilde{\tau}-\bar{x}_{\rm c}\cosh{\tilde{\eta}}\right)^{2}T^{\tilde{\tau}\tilde{\tau}}\\ &+\left(\bar{x}_{\rm c}\sinh{\tilde{\eta}}\right)^{2}\tilde{\tau}^{2}T^{\tilde{\eta}\tilde{\eta}}-\bar{x}_{\rm c}\sinh{\tilde{\eta}}\left(\tilde{\tau}-\bar{x}_{\rm c}\cosh{\tilde{\eta}}\right)\tilde{\tau}T^{\tilde{\tau}\tilde{\eta}}\Big{]}\ ,\tau^{2}T^{\eta\eta}\end{split}$ (36)
$\begin{split}&=\frac{1}{\tilde{\tau}^{2}-2\tilde{\tau}\bar{x}_{\rm c}\cosh{\tilde{\eta}}+\bar{x}^{2}_{\rm c}}\Big{[}\left(\bar{x}_{\rm c}\sinh{\tilde{\eta}}\right)^{2}T^{\tilde{\tau}\tilde{\tau}}\\ &+\left(\tilde{\tau}-\bar{x}_{\rm c}\cosh{\tilde{\eta}}\right)^{2}\tilde{\tau}^{2}T^{\tilde{\eta}\tilde{\eta}}-\bar{x}_{\rm c}\sinh{\tilde{\eta}}\left(\tilde{\tau}-\bar{x}_{\rm c}\cosh{\tilde{\eta}}\right)\tilde{\tau}T^{\tilde{\tau}\tilde{\eta}}\Big{]}\ ,\end{split}$ (37)
where $\bar{x}_{\rm c}=\sqrt{2}x_{\rm c}$. To estimate the energy–momentum tensor from lattice simulations, we use the definition of the energy–momentum tensor on a lattice given in our previous paper 3Dglasma6 .
3.1 Model Parameters
The set of model parameters for the initial $3$D color charge density includes the following: the longitudinal extension of the classical color charge density, $r_{L}$, the longitudinal correlation length in the classical color charge density, $l_{L}$, the parameter controlling the averaged square of the color charge density, $g^{2}\bar{\mu}$, and the infrared and ultraviolet cutoffs, $\Lambda_{\rm IR}$ and $\Lambda_{\rm UV}$. To simulate the Au–Au collision with a center-of-mass energy per nucleon pair of $200$ GeV at RHIC, we adjust $\sqrt{2}r_{L}$ to roughly match the radius of a gold nucleus divided by the gamma factor $\gamma=108$ and the factor $\sqrt{2}$ stemming from the definition of the light-cone variables, $r_{L}=0.1/Q_{\rm s}$, where the saturation scale is set at $Q_{\rm s}=1$ GeV. The parameter $g^{2}\bar{\mu}$ is set to ensure that the integration of the averaged square of the color charge density over the light-cone variable coincides with the squared saturation scale, $\int z\left(g^{2}\mu(x,\bm{x}_{\perp})\right)^{2}=Q^{2}_{s}$. This setup for $g^{2}\bar{\mu}$ indicates that the event average of a given observable has a translational invariance in the transverse plane. Such a translational invariance is approximately established in the central region of large nuclei in the transverse plane, where the thickness of the nucleus does not vary largely. Therefore, our setup simulates the glasma in the central region, as depicted in Figure 1. The infrared cutoff is set at $\Lambda_{\rm IR}=0.2Q_{\rm s}$, and the ultraviolet cutoff is taken to be double the value of $Q_{\rm s}$, $\Lambda_{\rm UV}=2Q_{\rm s}$. It is important to note that in the classical approximation the choice of the coupling constant only provides the trivial change in the field variables, as $A,E,J\propto 1/g$. This is because the coupling constant can be excluded in both the classical equations of motion and the continuity equation by redefining the scaled fields, $A^{\prime}=gA,E^{\prime}=gE$, and $J^{\prime}=gJ$. The numerical results shown in the later sections are normalized by $g$ to make them independent of $g$.
@@ 3.1 Model Parameters
We explore the effect of longitudinal correlation by comparing the 3+1D glasma under various correlation lengths, $l_{\rm L}$, ranging from $0.01r_{\rm L}$ to $2r_{\rm L}$. The smallest value, $l_{\rm L}=0.01r_{\rm L}$, which is significantly less than the nucleus radius, is thought to arise from the intricate structure within a nucleon. The largest value, $l_{\rm L}=2r_{\rm L}$, which is twice the nucleus radius, is the maximum limit discussed in Section 2.3. It is crucial to emphasize that the physical origin of the longitudinal correlation is beyond the scope of this paper.
In addition, the following parameters should be inputted for the 3D glasma simulation. These parameters are chosen to be sufficiently safe values to minimize their impact on the physical results. The transverse lattice grid number and lattice spacing are set to $N_{\perp}=56$ and $a_{\perp}=0.4/Q_{s}$, respectively. For the lattice in the modified rapidity direction $\tilde{\eta}$, the grid number and spacing are $L_{\tilde{\eta}}=896,7168$ and $a_{\tilde{\eta}}=2.6/L_{\tilde{\eta}}$, respectively. It is noteworthy that simulations with smaller $l_{\rm L}$ require smaller longitudinal lattice spacings to accurately capture the structure of the color charge density in the longitudinal direction. The beginning modified proper time in the simulation and the center positions of the nuclei are taken as $\tilde{\tau}_{\rm ini}=32r_{\rm L}$ and $x_{\rm c}=20\sqrt{2}r_{\rm L}$, respectively. All results in this section are estimated using $48$ events. The error in the results is estimated by dividing the unbiased variance by the square root of the number of events. According to the central limit theorem, the unbiased variance aligns with the variance of the mean value in the infinitely large number of events.
@@ 3.2 Midrapidity Region
First, we investigate the behavior of pressure and energy density in the mid-rapidity region, $\eta=0$. Figure 2 illustrates the evolution of glasma pressure at $\eta=0$, with the pressure normalized to the energy density. The time range $0\leq Q_{\rm s}\tau\leq 9.7$ covers the time before the onset time of the hydrodynamics, $\tau=0.6\sim 1.0$ Heinz , as predicted by the experimental analysis. The blue and red lines indicate pressures resulting from collisions of nuclei with two different lengths of longitudinal correlation in color charge densities, $l_{\rm L}=0.01r_{\rm L}$ and $2r_{\rm L}$. The solid and dotted lines correspond to transverse and longitudinal pressures over the energy density, $P_{\perp}/\varepsilon$ and $P_{\rm L}/\varepsilon$, respectively.
@@
Despite exploring two distinct longitudinal correlations in the color charge densities, the observed pressures were qualitatively similar, though they showed quantitative differences, particularly in the early time region. Notably, at $Q_{\rm s}\tau=0$, transverse pressure is minimal compared to longitudinal pressure, reflecting the overlap of two nuclei centers at $(\tau,\eta)=(0,0)$ and the dominance of the Weizsäcker–Williams field, a transverse electromagnetic field representing nuclei, which significantly contributes to longitudinal but not transverse pressure. Between $Q_{\rm s}\tau=0$ and $0.6$, both transverse and longitudinal pressures undergo substantial changes, eventually acquiring opposite signs. The realization of the relation $P_{\perp}>0>P_{\rm L}$ indicates the formation of a strong longitudinal color electromagnetic field between separating nuclei, like the boost invariant glasma. From $Q_{\rm s}\tau=0.6$ to $2$, the transverse pressure trends towards $0.5$, while longitudinal pressure approaches zero, with no significant changes observed beyond $\tau=2$. These findings are in qualitative agreement with the previous results from the glasma calculations, where other effects of boost invariance breaking are taken into account Instability14 ; Instability15 ; Instability16 ; Instability17 ; Instability18 ; Instability19 ; Instability20 ; Instability21 ; 3DJIMWLK .
Figure 3 illustrates the evolution of the glasma energy density $\varepsilon$ at $\eta=0$, normalized by that for $Q_{\rm s}\tau=1$, $\varepsilon_{0}$, where the solid blue and red lines depict the energy density arising from collisions of nuclei with two different lengths of longitudinal correlation in color charge densities, $l_{\rm L}=0.01r_{\rm L}$ and $2r_{\rm L}$. The time range is taken to be the same as in Figure 2. As in the pressure isotropization in Figure 2, despite the choice of $l_{\rm L}$, the results of the $3$D glasma simulation are the same qualitatively at late time. Between $Q_{\rm s}\tau=0$ and $0.6$, the energy density is dominated by the Weizsäcker–Williams field rather than the glasma. After $Q_{\rm s}\tau=0.6$, as the nuclei move past $\eta=0$ the evolution of the glasma’s energy density becomes apparent. A power function fitting with fitting parameters $A$ and $B$, $f_{\rm fit}(\tau)=A/\tau^{B}$, was employed to accurately determine the decay rate of the glasma energy density in the later stages. The obtained fitting parameter $B$, $B=1.06\pm 0.001$ for $l_{\rm L}=0.01r_{\rm L}$ and $B=1.04\pm 0.0004$ for $l_{\rm L}=2r_{\rm L}$, indicates that the energy of the glasma at $\eta=0$ decreases at almost the inverse of the proper time, $1/\tau$. This $1/\tau$ decrease in the energy density at late time has been observed by previous $2$D and $3$D glasma calculations Instability14 ; Instability16 ; Instability22 .
@@
The negligible longitudinal pressure, $P_{\perp}\gg P_{\rm L}$, and the decay of energy density, $\epsilon\propto 1/\tau$, in the late stage, as observed above, are consistent with the system expanding longitudinally like a dilute gas. These observations significantly deviate from the behavior expected in an ideal gas, where pressure is isotropic, $P_{\perp}=P_{\rm L}=P$, and the energy density decay is predicted by Bjorken hydrodynamics using the ideal gas equation of state, $\varepsilon=3P$, $\varepsilon\propto 1/\tau^{4/3}$.